Exhibit 9, Bushels demanded per month 50 Price per bushel S5 60 65 70 Bushels supplied per month 80 75 70 65 42-Which ot the following would occur if the government set a price ceiling of S1 in the market shown in Exhibit 9? a. There would be a shortage of apricots. b. Buyers would not want to purchase all of the apricots that are supplied. c. There would be a surplus of apricots d. Farmers would reduce the number of acres allocated to the growing of apricots 43-Which of the following would occur if the government sets a price floor of S4 in the market shown in Exhibit 97 a. There would be a shortage of apricots. b. Buyers would not purchase all of the apricots that are grown c. Buyers would purchase more apricots than are currently being supplied. d. Farmers would reduce the number of acres allocated to the growing of apricots 44- Suppose that during a recent year for the United States, merchandise imports were S2 trillion, unilateral transfers were a net outflow of S0.2 trillion, service exports were $0.2 trillion, service imports were S0.I trillio trillion. n, and merchandise exports were $1 .4 a. b. c. What was the merchandise trade deficit? What was the balance on goods and services? What was the current account balance? 45-A 5 percent increase in the price of digital apps reduces the amount of tablet devices demanded by 3 percent What i cross price elasticity of demand? Are tablet devices and digital apps complements or substitutes? below, characterize the demands for the following goods as being more elastic or 46- Based solely on the information provided more inelastic. the de a. A 45-cent box of salt that you buy once a year b. A type of high-powered ski boat th any one of aumber of rental agencies d. Automobile insurance in a state that requires e. A 75-cent guitar pick for the lead guitarist of a major rock band autos to be insured but has only a few insurance companies 47-. Scarcity a. is a problem only in the poorer countries of the world. b. can be solved by rapid advances in technology c. is a problem that exists in every economy d is not a problem for the very rich
Answer #1

42). At a price ceiling of $1, Qd is 70 and Qs is 55 so there is a shortage of 15 units. (Option A)

43). At a price floor of $4, Qd is 55 and Qs is 75 so there is a surplus of 20 units. (Option B)

44). Merchandise trade deficit = 1.4 - 2 = -0.6 trillion. Balance on goods and services = -0.6 + (0.2 - 0.1) = -0.5 trillion. CA balance = -0.5 + (-0.2) = -0.7 trillion

45). CPE = -3%/5% = -0.6. Since cross elasticity is negative the two goods are complements

46). Inelastic. Elastic, Elastic, Inelastic, Inelastic

47) Option C.